### v1

# Tokenizing Electron Cloud in Protein-Ligand Interaction Learning

Abstract
The affinity and specificity of protein-molecule binding directly impact functional outcomes, uncovering the mechanisms underlying biological regulation and signal transduction. Most deep-learning-based prediction approaches focus on structures of atoms or fragments. However, quantum chemical properties, such as electronic structures, are the key to unveiling interaction patterns but remain largely underexplored. To bridge this gap, we propose ECBind, a method for tokenizing electron cloud signals into quantized embeddings, enabling their integration into downstream tasks such as binding affinity prediction. By incorporating electron densities, ECBind helps uncover binding modes that cannot be fully represented by atom-level models. Specifically, to remove the redundancy inherent in electron cloud signals, a structure-aware transformer and hierarchical codebooks encode 3D binding sites enriched with electron structures into tokens. These tokenized codes are then used for specific tasks with labels. To extend its applicability to a wider range of scenarios, we utilize knowledge distillation to develop an electron-cloud-agnostic prediction model. Experimentally, ECBind demonstrates state-of-the-art performance across multiple tasks, achieving improvements of 6.42% and 15.58% in per-structure Pearson and Spearman correlation coefficients, respectively.
1 Introduction
Protein-ligand binding interaction is important for understanding the mechanisms of biological regulation, providing a theoretical basis for the design and discovery of new drugs (Fu et al., 2018). The binding process is characterized by two key attributes: (i) specificity, which enables the precise recognition of a highly specific binding partner over less specific alternatives; and (ii) affinity, which ensures that even a high concentration of weakly interacting partners cannot substitute for the effects of a low concentration of the specific partner with strong binding affinity (Du et al., 2016b). In the field of artificial-intelligence-aided drug design (AIDD), the learning of interaction patterns thus can be summarized into two tasks: (i) binary classification of active and inactive states of the binding complexes, and (ii) regression prediction of binding affinity.
In recent years, the abundance of protein-molecule binding structure data, coupled with the rise of 3D graph neural networks (GNNs), has significantly accelerated progress in the tasks (Satorras et al., 2022; Gilmer et al., 2017; Gasteiger et al., 2022). Most of these approaches (Gasteiger et al., 2024; Schütt et al., 2017; Stärk et al., 2022) treat the atoms or amino acids as fundamental units, modeling the binding graph at different scales (e.g., atom-level or fragment-level). By constructing translational, rotational, and reflectional equivariant or invariant features, they leverage the neural networks to extract geometric representations effectively and show promising performance in protein-ligand interaction tasks (Satorras et al., 2022; Geiger and Smidt, 2022).
One limitation is that their finest granularity is typically at the atomic level. However, at the quantum scale, electronic structures, compared to atoms within a molecule, reflect the fundamental forces governing intermolecular interactions (Leckband and Israelachvili, 2001). Specifically, beyond the 3D atomic structures represented by the ball-and-stick descriptions, which struggle to accurately capture interactions such as electrostatic forces, van der Waals interactions, hydrogen bonding, and $\pi$-$\pi$ stacking, electron cloud provides a detailed spatial distribution of electrons. It reveals critical features, such as charge polarization and lone pairs which are essential for understanding interaction patterns like non-covalent interactions and molecular polarity (Saleh et al., 2012). Moreover, in identifying binding sites, the shape and polarity of the electron cloud in molecules play a significant role in determining their affinity and selectivity for targets, as the polar regions of a drug molecule often complement those of the receptor, enhancing binding specificity (Isert et al., 2024). However, while perceiving electron clouds is crucial, integrating such descriptions of electronic structures into computational models poses significant challenges. This is due to (i) their inherent redundancy as natural signals and (ii) difficulty accessibility resulting from costly computational complexity (Schuch and Verstraete, 2009).
To bridge the gap, we introduce ECBind, a pretraining and finetuning paradigm designed to integrate protein-ligand electronic and atomic structures into an interaction prediction model. The pretraining phase enables the model to perceive both electron cloud as electronic structural descriptors and the full-atom structures of binding proteins and ligands, while the finetuning process allows the model to focus on specific tasks. Besides, to avoid the heavy computational cost in computing electron cloud, we develop a lightweight electron-cloud-agnostic student model with knowledge distillation. Empirically, ECBind demonstrates state-of-the-art performance across multiple tasks, achieving improvements of 6.42% and 15.58%16in per-structure Pearson and Spearman correlation coefficients on relative binding affinity prediction, with little performance degeneration in the student model.
2 Background
3D atomic binding graph. We define the binding system composed of a protein-ligand pair as $\mathcal{B}=(\bm{V},\bm{E})=(\mathcal{P},\mathcal{M})$, in which $\mathcal{P}$ is the protein and $\mathcal{M}$ is the molecule ligand. The node set of $\mathcal{P}$ is $\bm{V}_{\mathrm{rec}}$, containing $N_{\mathrm{rec}}$ atoms of proteins and $\bm{V}_{\mathrm{lig}}$ contains $N_{\mathrm{lig}}$ atoms of molecules. We denote the index set of the protein’s atom as $\mathcal{I}_{\mathrm{rec}}$ and the molecule’s atoms as $\mathcal{I}_{\mathrm{lig}}$, with $N_{\mathrm{fa}}=N_{\mathrm{rec}}+N_{\mathrm{lig}}$, leading to $\bm{V}_{\mathrm{rec}}=\{(a_{i},\bm{\mathrm{x}}^{\mathsf{a}}_{i})\}_{i\in\mathcal{I}_{\mathrm{rec}}}$ and $\bm{V}_{\mathrm{lig}}=\{(a_{i},\bm{\mathrm{x}}^{\mathsf{a}}_{i})\}_{i\in\mathcal{I}_{\mathrm{lig}}}$. Here, $a_{i}$ is the atom types with $a_{i}=1,\ldots,M$, and $\bm{\mathrm{x}}^{\mathsf{a}}_{i}\in\mathbb{R}^{3}$ is the corresponding 3D position. $\bm{E}$ is the edge set.
Electron cloud density. By employing numerical methods such as density functional theory (Capelle, 2006) or semi-empirical methods (Bannwarth et al., 2018), we can extract electron clouds from full atomic structures (including hydrogen atoms) as density maps of electronic structures. Typically, electron cloud is represented as $\mathcal{D}=\{(u_{i},\bm{\mathrm{x}}^{\mathsf{u}}_{i})\}_{1\leq i\leq N_{\mathrm{ec}}}$, where $u_{i}$ represents the electron density at position $\bm{\mathrm{x}}^{\mathsf{u}}_{i}$. Since the electron cloud signals are usually distributed on 3D grids, high-resolution signals inevitably result in $N_{\text{ec}}\gg N_{\text{fa}}$. Moreover, while the dense electron cloud contains sufficient interaction patterns compared to atom-level signals, it is excessively redundant. These lead to the following challenges:
  • From a computational perspective, if we treat $|\mathcal{D}|=N_{\text{ec}}$ as the number of nodes in the graph, message passing becomes infeasible due to the overwhelming computational cost. The reason is that usually, the scales of a binding site are about $\sim 20$Å, and if we set the resolution as 0.5 Å, it will leads $(20/0.5)^{3}$ nodes.).
  • From a signal processing perspective, such redundant signals carry limited semantic information and require further compression. Through compression can the resulting high-dimensional representations generalize more effectively and robustly to downstream tasks.
Interaction prediction. For the two attributes introduced in Sec. 1, we define the specificity prediction as finding a mapping $f$ such that $\hat{y}=f(\mathcal{B},\mathcal{D})\in[0,1]$, representing the probability of the ligand molecule’s activeness on the protein. Additionally, for the absolute binding affinity prediction, $\hat{y}_{p}=f(\mathcal{B}_{p},\mathcal{D}_{p})\in\mathbb{R}^{+}$ represents the binding strength in positive real values. For the relative binding affinity prediction, $\hat{y}_{pq}=f(\mathcal{B}_{p},\mathcal{D}_{p},\mathcal{B}_{q},\mathcal{D}_{q})\in\mathbb{R}$ indicates the difference in binding affinity as $\mathcal{B}_{q}=(\mathcal{P}_{q},\mathcal{M}_{q})$ changes to $\mathcal{B}_{p}=(\mathcal{P}_{q},\mathcal{M}_{p})$, with the corresponding changes in the electron cloud from $\mathcal{D}_{q}$ to $\mathcal{D}_{p}$. The encoder and decoder are all transformer-based, which can be regarded as graph neural networks for fully connected graphs.
3 Method
3.1 Tokenizing Electron Cloud
In order to (i) enable the learned embeddings to capture the electronic structure of the protein-ligand binding, (ii) reduce the redundancy in electron cloud signals generated by numerical methods, and (iii) enhance the generalizability of the learned representations for downstream tasks, we propose to pretrain an encoder with vector quantization as signal compression techniques to generate tokenized embeddings of the binding graph. Figure. 1 gives the workflows.
3.1.1 Encoder architecture
Figure 1: Pretraining electron-cloud-aware codebooks and finetuning on the specific tasks.
Patchifying electron clouds.
We first use linear projectors to lift $u_{i}$ into a higher-dimensional space, denoted as $\bm{z}^{\mathsf{u}}_{i}$, and similarly, $a_{i}$ is lifted as $\bm{z}^{\mathsf{a}}_{i}$. To downsample the resolution of the signal $\{u_{i}\}_{1\leq i\leq N_{\mathrm{ec}}}$, we patchify the electron cloud signals, where each patch contains $K$ electron densities, with its center being the nearest atom position. As a result, we obtain $N_{\mathrm{fa}}$ patches, whose embedding sequence is represented as $\bm{Z}^{\mathsf{u}}_{i}=(\bm{z}^{\mathsf{a}}_{i},\bm{z}^{\mathsf{u}}_{j_{1}},\ldots,\bm{z}^{\mathsf{u}}_{j_{K}})$, and the corresponding position sequence as $\bm{X}^{\mathsf{u}}_{i}=(\bm{\mathrm{x}}^{\mathsf{a}}_{i},\bm{\mathrm{x}}^{\mathsf{u}}_{j_{1}},\ldots,\bm{\mathrm{x}}^{\mathsf{u}}_{j_{K}})$. Here, ${j_{k}}$ is the subsequence of $(j)_{j=1,\ldots,N_{\mathrm{ec}}}$, such that $j_{k}$ lies in the neighborhood of $i$, determined by the Euclidean distance $||\bm{\mathrm{x}}^{\mathsf{u}}_{j_{k}}-\bm{\mathrm{x}}^{\mathsf{a}}_{i}||_{2}$. However, in practice, $K$ cannot be set as a sufficiently large value to cover the entire electron cloud signal set. To address it, we employ two techniques to allow the model to perceive electron densities farther from the patch center as part of the input:
  • Random Selection as Input Augmentation. Inspired by neural operators for spatially irregular inputs (Lin et al., 2021; Li et al., 2020a), we randomly sample points from the input grid, making the signals irregularly distributed to serve as inputs. Specifically, we use $r_{\mathrm{in}}$ to determine whether a given density value is included as part of the input.
  • Random Selection of Patch Elements. Another technique involves setting a large $K_{\max}$ for the neighborhood of each $i$-th atom and then randomly selecting $K$ electron cloud signals from the $K_{\max}$ neighbors. The selection is performed with a probability $p_{j_{k}}=\frac{\exp(\|\bm{\mathrm{x}}^{\mathsf{u}}_{j_{k}}-\bm{\mathrm{x}}^{\mathsf{a}}_{i}\|_{2}/\gamma_{\mathrm{pt}})}{\sum_{k=1}^{K_{\max}}\exp(\|\bm{\mathrm{x}}^{\mathsf{u}}_{j_{k}}-\bm{\mathrm{x}}^{\mathsf{a}}_{i}\|_{2}/\gamma_{\mathrm{pt}})}$, where $\gamma_{\mathrm{pt}}$ is the predefined smoothing coefficient, set as 0.1.
Structure-aware transformer. We refer to GeoMHA from ESM3 (Hayes et al., 2024) as the structure-aware module to construct the encoder’s network architecture. However, while ESM3 operates at the residue level, we adapt it to the atom/electron cloud level. Specifically, for $\bm{Z}\in\mathbb{R}^{L\times D}$ representing the embedding of one patch with length $L$, and $\bm{X}\in\mathbb{R}^{L\times 3}$ denoting the corresponding positions (where $D$ is the embedding dimension), a linear layer is first applied to project $\bm{Z}$ into invariant-feature keys and queries, geometric-feature keys and queries, and values, denoted as $\bm{Q}^{\mathsf{iv}},\bm{Q}^{\mathsf{ge}},\bm{K}^{\mathsf{iv}},\bm{K}^{\mathsf{ge}},\bm{V}\in\mathbb{R}^{L\times H\times 3}$, where $H$ is the number of attention heads. The first attention matrix, $\bm{\mathrm{A}}^{\mathsf{iv}}\in\mathbb{R}^{L\times L\times H}$, regulates the classical attention mechanism, while the second, $\bm{\mathrm{A}}^{\mathsf{ge}}\in\mathbb{R}^{L\times L\times H}$, incorporates geometric-awareness. Their difference collectively forms the final attention weights, $\bm{\mathrm{A}}$, enabling message-passing among $L$ input points, given by:
$\bm{\mathrm{A}}$ $=\mathrm{softmax}(\bm{\mathrm{A}}^{\mathsf{iv}}-\bm{\mathrm{A}}^{\mathsf{ge}})$ (1)
$\text{where}\quad\bm{\mathrm{A}}^{\mathsf{iv}}=\frac{\mathrm{softplus}(w^{\mathsf{iv}})}{\sqrt{3}}(\bm{Q}^{\mathsf{iv}})^{\intercal}\bm{K}^{\mathsf{iv}},\quad\text{and}\quad\bm{\mathrm{A}}^{\mathsf{ge}}=\frac{\mathrm{softplus}(w^{\mathsf{ge}})}{\sqrt{3}}\mathrm{pdist}(\bm{X}\cdot\bm{Q}^{\mathsf{ge}},\bm{X}\cdot\bm{K}^{\mathsf{ge}}).$
$w^{\mathsf{iv}}$ and $w^{\mathsf{ge}}$ are learnable scalars, and $\mathrm{pdist}$ is the pairwise-distance function. The output is computed as $\bm{O}=\bm{\mathrm{A}}\bm{V}$. Finally, a feed-forward network (FFN) with LayerNorm and residual connections projects $\bm{O}$ back into the space of $\bm{Z}$. Appendix. B gives a brief proof of the trans-rotational invariance of the extracted embedding $\bm{Z}$.
In specific, for a patch defined by $\bm{Z}^{\mathsf{u}}_{i}$ and $\bm{X}^{\mathsf{u}}_{i}$, the patch length is given by $L=K+1$. After stacking multiple GeoMHA-based transformer layers to form the encoder $\phi^{\mathsf{ec}}(\cdot,\cdot)$, the representation of the patch center (indexed by 1) is extracted as the final representation of the patch, denoted as $\bm{h}^{\mathsf{u}}_{i}=[\phi^{\mathsf{ec}}(\bm{Z}^{\mathsf{u}}_{i},\bm{X}^{\mathsf{u}}_{i})]_{1}$.
3.1.2 Codebook construction
The patchwise representation of $\bm{h}^{\mathsf{u}}_{i}$ will then be tokenized into quantized codes. Following recent advances in language modeling as different modalities require specific prompts, hierarchical codebook is used and decomposed into two types: 3D electronic structures and 2D atomic attributes, denoted as $\mathcal{A}^{\mathsf{u}}=\mathcal{A}^{\mathsf{u}(1)}\times\mathcal{A}^{\mathsf{u}(2)}$. Each $\mathcal{A}^{\mathsf{u}(k)}$ is parameterized with learnable embeddings $\{\bm{e}^{\mathsf{u}(k)}_{j}\}_{j=1}^{|\mathcal{A}^{\mathsf{u}(k)}|}$.
To generate the tokenized codes for different levels, $\bm{h}^{\mathsf{u}}_{i}$ is first split as $\bm{h}^{\mathsf{u}}_{i}=[\bm{h}^{\mathsf{u}(1)}_{i};\bm{h}^{\mathsf{u}(2)}_{i}]$. Then, $\bm{h}^{\mathsf{u}(k)}_{i}$ is tokenized into discrete codes $\bm{h}_{i}^{\prime\mathsf{u}(k)}$ using vector quantization, as
$\bm{h}_{i}^{\prime\mathsf{u}(k)}=\arg\min_{j}\|\bm{h}_{i}^{\mathsf{u}(k)}-\bm{e}_{j}^{\mathsf{u}(k)}\|_{2},\quad k=1,2.$ (2)
In practice, we employ spherical projection to regularize $\bm{h}_{i}^{\mathsf{u}(k)}$ and $\bm{e}_{j}^{\mathsf{u}(k)}$, which enhances training stability (Huh et al., 2023). Additionally, Repeated K-means (Chiu et al., 2022) is applied to mitigate codebook collapse.
3.1.3 Structure reconstruction
Electron cloud reconstruction. By compressing $N_{\mathrm{ec}}$ signals into $N_{\mathrm{fa}}$ tokens, we aim for $\bm{H}^{\prime\mathsf{u}(1)}=\{\bm{h}^{\prime\mathsf{u}(1)}_{i}\}_{1\leq i\leq N_{\mathrm{fa}}}$ to retain key electronic structure information. Therefore, to pretrain the tokenizer composed of encoder $\phi^{\mathsf{ec}}$ and codebook $\mathcal{A}^{\mathsf{u}(1)}$, we hope to reconstruct the $N_{\mathrm{ec}}$ electron cloud signals using only $N_{\mathrm{fa}}$ patch-level embeddings. In detail, we first use layers of classical transformers for patch-level message passing, as $\bm{H}^{\mathsf{pt}(1)}=\psi^{\mathsf{pt}}(\bm{H}^{\prime\mathsf{u}(1)})$. Then the patch-to-electron message-passing can be performed similarly to the structure-aware transformer in Sec. 3.1.1, with
$\quad\quad\quad\bm{h}_{j}^{\mathsf{pe}}$ $=[\psi^{\mathsf{pe}}(\bm{H}^{\mathsf{pe}(1)}_{j},\bm{{X}}^{\mathsf{pe}}_{j})]_{1},\quad\quad$ (3)
$,\displaystyle\text{where}\quad\bm{H}^{\mathsf{pe}(1)}_{j}=(\bm{0},\bm{h}_{i_{1}}^{\mathsf{pt}(1)},$$\ldots,\bm{h}_{i_{K}}^{\mathsf{pt}(1)}),\quad\quad\text{ and}\quad\bm{X}^{\mathsf{pe}}_{j}=(\bm{\mathrm{x}}^{\mathsf{u}}_{j},\bm{\mathrm{x}}_{i_{1}}^{\mathsf{pt}},\ldots,\bm{\mathrm{x}}_{i_{K}}^{\mathsf{pt}}).$
$\bm{0}$ is a zero vector like [CLS] in language modeling, $(i_{k})_{1\leq k\leq K}$ is the subsequence of $(i)_{1\leq i\leq N_{\mathrm{fa}}}$ and forms the $K$-nearest neighborhood of $j$ according to their distance. The patch-to-electron module will generate $N_{\mathrm{ec}}$ embeddings of $\{\bm{h}^{\mathsf{pe}}_{j}\}_{1\leq j\leq N_{\mathrm{ec}}}$. And the final embeddings of the electron cloud will be computed with another classical transformer $\bm{H}^{\mathsf{ec}}=\psi^{\mathsf{ec}}(\{\bm{h}^{\mathsf{pe}}_{j}\}_{1\leq j\leq N_{\mathrm{ec}}})\in\mathbb{R}^{N_{\mathrm{ec}}\times D}$, allowing electron-level message-passing. A linear projector is used to map $\bm{h}_{j}^{\mathsf{ec}}$ to $\hat{u}_{j}$ and Mean Square Error (MSE) is employed to reconstruct the electron clouds,
$L^{\mathsf{ec}}=\mathrm{MSE}_{j}(\hat{u}_{j},u_{j}),\quad\quad\text{where }1\leq j\leq N_{\mathrm{ec}}.$ (4)
$\mathrm{MSE}_{j}(\cdot,\cdot)$ indicates mean reduction across the index $j$.
Masked atom type prediction. Moreover, to enable the encoder to perceive the 2D graph topology, we use $\bm{h}^{\prime\mathsf{u}(2)}_{i}$, obtained from the codebook $\mathcal{A}^{\mathsf{u}(2)}$, as the embedding of 2D atomic attributes. Since each patch center is defined as the nearest atom position, $\bm{h}^{\prime\mathsf{u}(2)}_{i}$ can naturally be regarded as the embedding of the $i$-th atom. To train the tokenized representation effectively, we adopt techniques from Masked AutoEncoder (MAE) (He et al., 2021). Specifically, a ratio of atom types in the input is randomly selected and masked (Masking Ratio is set as 0.1), and the tokenizer extracts $\bm{h}^{\prime\mathsf{u}(2)}_{i}$ from the randomly masked inputs. Subsequently, an atom-level transformer is applied to decode $\bm{h}^{\prime\mathsf{u}(2)}_{i}$ into $\mathrm{logit}_{i}^{\mathsf{at}}\in\mathbb{R}^{M}$, as the predicted logits of the masked atom types’ probabilities with cross-entropy (CE) loss utilized as
$L^{\mathsf{at}}=\text{CE}_{i}(\mathrm{logit}_{i}^{\mathsf{at}},a_{i}),\quad 1\leq i\leq N_{\mathrm{fa}}\text{ and }a_{i}\text{ is masked}.$ (5)
Since the encoder $\phi^{\mathsf{ec}}$ are shared by the two tasks, we can regard this prediction task as using the contextual of unmasked atom types with the positions and the electron cloud shape to predict the masked atom types correctly.
3.2 Tokenizing Atom-level Structure
The electronic structure usually provides sophisticated information for local interaction, and besides it, we aim for our model to perceive a more global structural one. To achieve this, we establish an atom-level tokenizer that encodes each atom with its macro-environment (Wu et al., 2024a), enabling the model to capture long-range inter-atomic interactions. We use a full-atom GeoMHA-based transformer to encode the macro-environment (See Appendix. C.1), and a classical transformer to decode the tokenized representation into a structural one (See Appendix. C.2). Two types of loss functions are used: (i) intra-molecular binned distance classification loss and (ii) inter-molecular interaction loss. (See Appendix. C.2)
3.3 Training, Usage, and Distillation
Training. We employ a pretrain-finetune paradigm to train our proposed ECBind. For the electron cloud tokenizer pretraining, besides the discussed loss, a commitment loss is also employed:
$L^{\mathsf{cmt}(k)}=\text{MSE}_{i}(\text{sg}(\bm{h}^{\mathsf{u}(k)}_{i}),\bm{e}_{i}^{\mathsf{u}(k)}),$ (6)
where the update for $\bm{e}_{j}^{\mathsf{u}(k)}$ is performed using an Exponential Moving Average (EMA), and $\text{sg}(\cdot)$ denotes the stop-gradient operation (Bengio et al., 2013). The training objective for the tokenizer is:
$L^{\mathsf{u}}=L^{\mathsf{ec}}+L^{\mathsf{at}}+\alpha^{\mathsf{u}}(L^{\mathsf{cmt}(1)}+L^{\mathsf{cmt}(2)}),$ (7)
where $\alpha^{\mathsf{u}}$ is the loss weight, typically set to 10.0. The atom-level tokenizer training follows the same.
Usage. For downstream tasks such as binding affinity prediction, we employ an attention mechanism to adaptively learn the combination of codebooks as the representation:
$\bm{h}_{i}=\sum_{\mathsf{s}}\alpha_{i}^{\mathsf{s}}$ $\bm{h}_{i}^{\prime}{{}^{\mathsf{s}}};\quad\quad\alpha_{i}^{\mathsf{s}}=\varphi(\{\bm{h}_{i}^{\prime}{{}^{\mathsf{s}}}\}),$ (8)
${\mathsf{s}\in\{\mathsf{u}(1),\mathsf{u}(2),\mathsf{a}(1),\mathsf{a}(2)\}}.$
A lightweight network $\psi(\cdot)$ is subsequently stacked to decode $\bm{H}=\{\bm{h}_{i}\}_{1\leq i\leq N_{\text{fa}}}$ into the target value with task-specific objectives. For example, for absolute binding affinity prediction, the $p$-th complex structure generates the binding affinity ${y}_{p}$ as ground truth, and $\hat{y}_{p}=\psi(\bm{H}_{p})\in\mathbb{R}$ is fine-tuned with the objective $L^{\mathsf{ba}}=\text{MSE}(\hat{y}_{p},y_{p})$ ; For relative affinity prediction, the label is ${y}_{pq}=y_{p}-y_{q}$, and $\hat{y}_{pq}=\psi([\bm{H}_{p},\bm{H}_{q}])$ is trained with $L^{\mathsf{ba}}=\text{MSE}_{i}(\hat{y}_{pq},y_{pq}).$
Distillation. However, obtaining electron cloud signals is computationally expensive. To address this, we employ a student network, sharing the same architecture as the trained teacher, to distill knowledge (Hinton et al., 2015; Zhang et al., 2020; Wu et al., 2022). In detail, the student approximates the merged tokens, avoiding the redundancy of electron cloud signals and reducing error accumulation. Instead of building an electron cloud approximator, we use only the full-atom structure of the binding site as input, with tokenized embeddings trained to approximate $\bm{h}_{i}$. Notably, we find that the electron-cloud-agnostic student network achieves comparable performance (see Sec. 5).
4 Related Work
Tokenizing biomolecules. The success of quantization techniques for image generation and representation (Yu et al., 2022; Esser et al., 2021; Lee et al., 2022) boosts works of tokenizing biomolecules. Foldseek is motivated by expanding the codebook of 20 amino acids for fast retrieving protein data (van Kempen et al., 2022). Most existing methods tokenize atoms or fragments (e.g., amino acids) with their surrounding environments, enabling models to perceive molecular structure and topology (Su et al., 2024; Yuan et al., 2023; Pourmirzaei et al., 2024; Fu et al., 2024; Li et al., 2024) for downstream tasks such as property prediction or structure generation (Wu et al., 2024a, b). Recently, ESM3 (Hayes et al., 2024), as a multimodal language model, generates tokens for reasoning over the sequence, structure, and function of proteins. Alternatively, methods focusing on finer-grained representations, such as surface point clouds extracted by chemical software or algorithms (Sverrisson et al., 2020), have also been proposed (Song et al., 2024; Wu and Li, 2024; Zhang et al., 2025). While biomolecular surfaces provide additional complements, they are not as natural and fundamental as electron clouds andd hardly reveal the underlying interactions between biomolecules.
Learning interaction. Interaction between biomolecules  (Du et al., 2016a) are widely explored in the machine learning community (Luo et al., 2023; Somnath et al., 2021; Wang et al., 2022), which are ususlly based on SE(3)-equivariance (Liao and Smidt, 2023; Liao et al., 2024; Aykent and Xia, 2025). For molecules as ligands, they are usually represented as graphs at full-atom level (Hoogeboom et al., 2022; Xu et al., 2022; Lin et al., 2022) or fragments (Geng et al., 2023; Jin et al., 2018; Lin et al., 2023a). Besides, hierarchical ones are also proposed, achieving the state-of-the-art performance (Kong et al., 2024; Cen et al., 2024). These models are established with mostly supervised training. For unsupervised ones (Jin et al., 2023b, a) including pretraining (Wu et al., 2024b; Wu and Li, 2024), they aim to enable the model to perceive structural flexibility. However, the electronic structure (Zhang et al., 2024; Adams et al., 2024) governing the binding modes is largely unexplored in binding affinity prediction.
Relation to neural operators. The proposed task of reconstruction of electron cloud density is similar to the motivation of neural operators (Li et al., 2021; Kovachki et al., 2023; Lu et al., 2021). Firstly, neural operators are well-suited for processing signals with varying resolutions. Under the irregularly spatially sampled setting, our random selection as input augmentation in Sec. 3.1.1 aligns closely with the strengths of graph kernel neural operators (Li et al., 2020a). Furthermore, the transformer with an attention mechanism is inherently a type of neural operator (Wright and Gonzalez, 2021), which employs non-Mercer kernel functions to adaptively learn message-passing patterns through iterative updates (Dosovitskiy et al., 2021; Guibas et al., 2022). Additionally, in the decoding process from patch centers to electron cloud signals, the coarse-to-fine message-passing resembles the multi-resolution matrix factorization employed in multipole graph neural operators (Li et al., 2020b). The effectiveness of such network architectures and training augmentations in neural operators has been extensively validated (Lin et al., 2023b; Pathak et al., 2022), and serves as one of the key inspirations for our design.
5 Experiments
Table 1: Results of relative binding affinity prediction on MISATO. Bold metrics are the best, and the underlined values are the best baseline except ECBind.
Methods Metrics Overall Per-structure
Pearson($\uparrow$) Spearman($\uparrow$) RMSE($\downarrow$) AUROC($\uparrow$) Pearson($\uparrow$) Spearman($\uparrow$)
Benchmark GCN 0.502$\pm$0.038 0.421$\pm$0.043 4.388$\pm$0.127 0.741$\pm$0.020 0.163$\pm$0.015 0.182$\pm$0.018
Fragment EGNN 0.579$\pm$0.037 0.535$\pm$0.012 4.572$\pm$0.124 0.759$\pm$0.019 0.297$\pm$0.029 0.244$\pm$0.038
SchNet 0.590$\pm$0.024 0.474$\pm$0.012 4.566$\pm$0.052 0.677$\pm$0.007 0.322$\pm$0.036 0.307$\pm$0.030
TorchMD 0.522$\pm$0.038 0.392$\pm$0.052 4.465$\pm$0.112 0.661$\pm$0.034 0.273$\pm$0.047 0.242$\pm$0.053
LEFTNet 0.639$\pm$0.043 0.497$\pm$0.059 4.385$\pm$0.091 0.717$\pm$0.019 0.284$\pm$0.024 0.267$\pm$0.021
GemNet 0.620$\pm$0.035 0.385$\pm$0.089 4.285$\pm$0.040 0.624$\pm$0.056 0.308$\pm$0.039 0.292$\pm$0.011
DimeNet 0.446$\pm$0.063 0.428$\pm$0.035 4.762$\pm$0.138 0.695$\pm$0.033 0.024$\pm$0.083 0.051$\pm$0.041
Equiformer 0.666$\pm$0.020 0.499$\pm$0.045 4.371$\pm$0.063 0.698$\pm$0.051 0.339$\pm$0.011 0.302$\pm$0.044
MACE 0.592$\pm$0.044 0.427$\pm$0.035 4.740$\pm$0.055 0.699$\pm$0.055 0.316$\pm$0.052 0.262$\pm$0.018
GET 0.651$\pm$0.030 0.473$\pm$0.005 4.494$\pm$0.056 0.673$\pm$0.009 0.366$\pm$0.013 0.359$\pm$0.009
HEGNN 0.555$\pm$0.083 0.497$\pm$0.081 4.467$\pm$0.112 0.697$\pm$0.054 0.262$\pm$0.011 0.235$\pm$0.002
Atom EGNN 0.378$\pm$0.099 0.321$\pm$0.048 4.938$\pm$0.068 0.548$\pm$0.015 0.248$\pm$0.005 0.242$\pm$0.042
SchNet 0.396$\pm$0.071 0.344$\pm$0.073 4.901$\pm$0.130 0.597$\pm$0.056 0.222$\pm$0.028 0.233$\pm$0.023
TorchMD 0.427$\pm$0.131 0.351$\pm$0.115 4.829$\pm$0.110 0.703$\pm$0.067 0.155$\pm$0.031 0.145$\pm$0.035
LEFTNet 0.364$\pm$0.098 0.270$\pm$0.037 4.881$\pm$0.050 0.517$\pm$0.019 0.196$\pm$0.025 0.173$\pm$0.057
GemNet 0.578$\pm$0.048 0.376$\pm$0.072 4.624$\pm$0.090 0.660$\pm$0.050 0.375$\pm$0.032 0.348$\pm$0.056
DimeNet 0.393$\pm$0.147 0.336$\pm$0.217 5.040$\pm$0.014 0.677$\pm$0.045 0.071$\pm$0.033 0.100$\pm$0.062
Equiformer OOM OOM OOM OOM OOM OOM
MACE 0.452$\pm$0.172 0.359$\pm$0.162 4.469$\pm$0.351 0.641$\pm$0.099 0.237$\pm$0.010 0.194$\pm$0.065
GET 0.443$\pm$0.079 0.307$\pm$0.089 4.913$\pm$0.080 0.582$\pm$0.007 0.294$\pm$0.048 0.267$\pm$0.030
HEGNN 0.545$\pm$0.089 0.486$\pm$0.069 4.501$\pm$0.077 0.705$\pm$0.049 0.267$\pm$0.027 0.271$\pm$0.039
Hierarchical EGNN 0.643$\pm$0.003 0.531$\pm$0.044 4.265$\pm$0.130 0.723$\pm$0.045 0.333$\pm$0.056 0.355$\pm$0.041
SchNet 0.601$\pm$0.098 0.520$\pm$0.033 4.328$\pm$0.137 0.709$\pm$0.013 0.350$\pm$0.048 0.336$\pm$0.065
TorchMD 0.597$\pm$0.029 0.420$\pm$0.034 4.253$\pm$0.131 0.670$\pm$0.031 0.329$\pm$0.015 0.316$\pm$0.031
LEFTNet 0.611$\pm$0.026 0.476$\pm$0.059 4.230$\pm$0.051 0.718$\pm$0.058 0.350$\pm$0.015 0.374$\pm$0.054
GemNet 0.655$\pm$0.067 0.547$\pm$0.080 4.135$\pm$0.113 0.769$\pm$0.068 0.248$\pm$0.108 0.244$\pm$0.122
DimeNet 0.599$\pm$0.020 0.509$\pm$0.059 4.607$\pm$0.049 0.699$\pm$0.026 0.346$\pm$0.016 0.293$\pm$0.098
Equiformer OOM OOM OOM OOM OOM OOM
MACE 0.608$\pm$0.097 0.549$\pm$0.048 4.035$\pm$0.464 0.783$\pm$0.052 0.240$\pm$0.114 0.206$\pm$0.084
GET 0.646$\pm$0.016 0.472$\pm$0.048 4.423$\pm$0.063 0.657$\pm$0.028 0.405$\pm$0.019 0.385$\pm$0.056
HEGNN 0.590$\pm$0.028 0.535$\pm$0.029 4.333$\pm$0.253 0.762$\pm$0.034 0.313$\pm$0.013 0.291$\pm$0.035
Ours ECBind-nptn 0.621$\pm$0.026 0.554$\pm$0.014 3.936$\pm$0.025 0.759$\pm$0.014 0.283$\pm$0.016 0.258$\pm$0.028
ECBind-stdt 0.697$\pm$0.019 0.646$\pm$0.008 3.907$\pm$0.122 0.824$\pm$0.021 0.407$\pm$0.027 0.420$\pm$0.025
ECBind-ptn 0.692$\pm$0.013 0.665$\pm$0.014 3.892$\pm$0.044 0.843$\pm$0.033 0.431$\pm$0.022 0.445$\pm$0.034
Figure 2: Visualization: GET (blue) and ECBind (red) on per-structure metrics for ‘hard’ protein samples.
5.1 Setup
Tasks and Datasets. We evaluate the effectiveness of the proposed ECBind on three tasks and their corresponding datasets. First, we follow the protocol from Atom3D (Townshend et al., 2020), which includes all protein amino acids within a distance of $\leq 6\ \text{\AA}$ from the atoms in the molecule to form the pocket. We account for the positions of hydrogen atoms, which is crucial for accurate electron cloud computation. Specifically, we use Reduce (Grosse-Kunstleve et al., 2002) to add missing hydrogen atoms in the proteins’ amino acids and compute their coordinates, while MMFF (Tosco et al., 2014) is employed to estimate the coordinates of missing hydrogens in the molecule. After obtaining the refined binding conformations, we use XTB (Bannwarth et al., 2018) to compute the electron cloud density. Structures that fail to converge during computation are treated as low-quality data and excluded from the datasets. This leads to the listed tasks and datasets:
  • (i)
    Relative binding affinity with MISATO (Siebenmorgen et al., 2024): It refines PDBBind (Wang et al., 2004) using molecular dynamics simulations of 19,443 experimental protein-ligand complexes. However, we exclude the quantum mechanical properties it provides, as they are challenging to obtain in real-world scenarios. After processes, including the removal of data without labels and filtering based on convergence during electron cloud computation, the final training, validation, and test sets are split according to the benchmark [^1] https://github.com/kierandidi/misato-affinity, resulting in 9,964, 1,074, and 156 pairs.
  • (ii)
    Absolute binding affinity with LBA (Liu et al., 2014): Using a sequence identity split of 30%, the included pairs are processed similarly to (i), resulting in 3,067, 346, and 426 instances for training, validation, and test.
  • (iii)
    Ligand specificity with LEP (Friesner et al., 2004): With split by functional proteins, the dataset yields 536, 202, and 162 pairs for training, validation, and test sets.
Table 2: Results of metrics on LBA and LEP. Since they are absolute binding prediction tasks, ‘per-structure’ metrics are not feasible. Bold metrics are the best, and the underlined values are the best baseline except ECBind.
Methods Metrics LBA LEP
Pearson($\uparrow$) Spearman($\uparrow$) RMSE($\downarrow$) AUROC($\uparrow$) AUPRC($\uparrow$)
Block SchNet 0.565$\pm$0.006 0.549$\pm$0.007 1.406$\pm$0.020 0.732$\pm$0.022 0.718$\pm$0.031
DimeNet++ 0.576$\pm$0.016 0.569$\pm$0.016 1.391$\pm$0.020 0.669$\pm$0.014 0.609$\pm$0.036
EGNN 0.566$\pm$0.010 0.548$\pm$0.012 1.409$\pm$0.015 0.746$\pm$0.017 0.755$\pm$0.031
ET 0.599$\pm$0.017 0.584$\pm$0.025 1.367$\pm$0.037 0.744$\pm$0.034 0.721$\pm$0.052
GemNet 0.569$\pm$0.027 0.553$\pm$0.026 1.393$\pm$0.036 0.590$\pm$0.062 0.541$\pm$0.107
MACE 0.599$\pm$0.010 0.580$\pm$0.014 1.385$\pm$0.006 0.617$\pm$0.022 0.549$\pm$0.007
Equiformer 0.604$\pm$0.013 0.591$\pm$0.012 1.350$\pm$0.019 0.000$\pm$0.000 0.000$\pm$0.000
LEFTNet 0.588$\pm$0.011 0.576$\pm$0.010 1.377$\pm$0.013 0.738$\pm$0.001 0.750$\pm$0.006
GET 0.611$\pm$0.005 0.610$\pm$0.010 1.356$\pm$0.004 0.766$\pm$0.010 0.717$\pm$0.033
HEGNN 0.589$\pm$0.013 0.607$\pm$0.011 1.422$\pm$0.023 0.739$\pm$0.003 0.710$\pm$0.038
Atom SchNet 0.598$\pm$0.011 0.592$\pm$0.015 1.357$\pm$0.017 0.712$\pm$0.026 0.639$\pm$0.033
DimeNet++ 0.547$\pm$0.015 0.536$\pm$0.016 1.439$\pm$0.036 0.589$\pm$0.049 0.503$\pm$0.020
EGNN 0.599$\pm$0.002 0.587$\pm$0.004 1.358$\pm$0.000 0.711$\pm$0.020 0.643$\pm$0.041
ET 0.591$\pm$0.007 0.583$\pm$0.009 1.381$\pm$0.013 0.677$\pm$0.004 0.636$\pm$0.054
GemNet 0.610$\pm$0.004 0.580$\pm$0.007 1.341$\pm$0.009 0.575$\pm$0.012 0.482$\pm$0.033
MACE 0.579$\pm$0.009 0.563$\pm$0.012 1.411$\pm$0.029 0.614$\pm$0.047 0.571$\pm$0.046
Equiformer OOM OOM OOM OOM OOM
LEFTNet 0.610$\pm$0.004 0.598$\pm$0.003 1.343$\pm$0.004 0.740$\pm$0.024 0.698$\pm$0.026
GET 0.613$\pm$0.010 0.618$\pm$0.009 1.309$\pm$0.014 0.769$\pm$0.027 0.736$\pm$0.034
HEGNN 0.591$\pm$0.007 0.596$\pm$0.006 1.521$\pm$0.017 0.725$\pm$0.056 0.638$\pm$0.099
Hierarchical SchNet 0.590$\pm$0.017 0.571$\pm$0.028 1.370$\pm$0.028 0.736$\pm$0.020 0.731$\pm$0.048
DimeNet++ 0.582$\pm$0.009 0.574$\pm$0.007 1.388$\pm$0.010 0.579$\pm$0.118 0.517$\pm$0.100
EGNN 0.586$\pm$0.004 0.568$\pm$0.004 1.380$\pm$0.015 0.724$\pm$0.027 0.720$\pm$0.056
ET 0.580$\pm$0.008 0.564$\pm$0.004 1.383$\pm$0.009 0.717$\pm$0.033 0.724$\pm$0.055
GemNet 0.535$\pm$0.041 0.535$\pm$0.016 1.455$\pm$0.081 0.546$\pm$0.023 0.519$\pm$0.070
MACE 0.612$\pm$0.010 0.592$\pm$0.010 1.372$\pm$0.021 0.530$\pm$0.032 0.456$\pm$0.038
Equiformer OOM OOM OOM OOM OOM
LEFTNet 0.592$\pm$0.014 0.580$\pm$0.011 1.366$\pm$0.016 0.719$\pm$0.020 0.708$\pm$0.012
GET 0.614$\pm$0.005 0.612$\pm$0.013 1.309$\pm$0.011 0.767$\pm$0.017 0.738$\pm$0.043
HEGNN 0.586$\pm$0.005 0.592$\pm$0.007 1.486$\pm$0.011 0.703$\pm$0.028 0.688$\pm$0.030
Ours ECBind-nptn 0.613$\pm$0.008 0.606$\pm$0.008 1.359$\pm$0.027 0.743$\pm$0.007 0.752$\pm$0.005
ECBind-stdt 0.614$\pm$0.004 0.615$\pm$0.003 1.314$\pm$0.005 0.768$\pm$0.006 0.777$\pm$0.000
ECBind-ptn 0.626$\pm$0.006 0.622$\pm$0.004 1.298$\pm$0.003 0.775$\pm$0.008 0.763$\pm$0.004
(a) Electron cloud tokenzier loss
(b) Full atom tokenizer loss
(c) Overall correlations
Figure 4. Losses in tokenizers when training and testing, and metrics of variants of ECBind with/without different tokenizers.
Training and Baseline. For performance comparison, we follow the protocols of GET (Kong et al., 2024), benchmarking against GNN baselines including SchNet (Schütt et al., 2017), DimNet++ (Gasteiger et al., 2022), GemNet (Gasteiger et al., 2024), TorchMD-Net (Thölke and Fabritiis, 2022), LEFTNet (weitao Du et al., 2023), MACE (Batatia et al., 2022), and Equiformer (Liao and Smidt, 2023), and the latest proposed HEGNN (Cen et al., 2024). The descriptions of these models can be found in Appendix D. Following the categorization from GET, we group the baselines into fragment-level, atom-level, and hierarchical message-passing methods. In ECBind, we pretrain both atomic and electronic structures and finetune the lightweight transformer head on the same dataset for specific tasks, leading to ECBind-ptn. For other baseline models, they are directly trained on the training sets, with hyperparameters detailed in Appendix E. Additionally, we introduce two variations of ECBind: ECBind-nptn, a non-pretraining version that only leverages full-atom structure as inputs, to demonstrate the performance gains from pretraining and the additional gaining from electron cloud information, and ECBind-stdt as an electron-cloud-agnostic student version trained with knowledge distillation from ECBind-ptn.
[…]
5.2 Results
MISATO relative binding affinity. Table. 1 gives the performance comparison of ECBind with different baseline models, and Figure. 2 gives the visualization in certain ‘hard’ proteins with different molecules that models hardly give accurate predictions, from which we can conclude that
  • (i)
    ECBind with pretrained tokens achieves the overall best according to the reported metrics, with 6.42% and 15.58% improvements on per-structure Pearson and Spearman correlation over the best baseline model, respectively.
  • (ii)
    The non-pretrained version shows comparable performance in overall metrics; However, it struggles to differentiate variations in affinity caused by different molecules binding to the same protein, as shown in its significantly lower per-structure Pearson and Spearman correlations. Additionally, the electron-cloud-agnostic student model exhibits a marginal performance drop compared to the teacher model, with 5.56% and 5.61% decreases in Pearson and Spearman correlation.
  • (iii)
    Overall, models incorporating fragment-level graphs (Fragment&Hierarchical) outperform those based solely on atom-level graphs. In comparison, ECBind integrates electron-cloud- and atom-level structures. This suggests a promising direction for our future work—incorporating amino acids and functional groups into ECBind to further enhance its capabilities.
LBA and LEP. Table. 2 shows the results of binding affinity and specificity prediction on LBA and LEP and demonstrates the superiority of ECBind in the two tasks. It is worth noting that while ECBind achieves the best performance in overall Pearson and Spearman correlations on the LBA, its improvement over the non-pretrained version and the previous state-of-the-art GET is not as significant on MISATO. This will be further discussed in the next section.
5.3 Analysis
Figure 5: Representation similarity of ‘key atoms’ from non-pretrained and pretrained ECBind.
To figure out what contributes to the performance gain, we suppose and analyze two factors in pretraining in detail. Besides, an ablation study is provided in Appendix. F on hyper-parameter sensitivity, and Appendix. G gives cost of inference time and training memory.
Accurate reconstruction of electron cloud helps. We hypothesize that the relatively modest performance gains on the LBA dataset may stem from factors related to dataset size and quality. On one hand, the MISATO dataset is over twice the size of the LBA dataset. On the other hand, MISATO benefits from noise reduction through quantum chemistry calculations, molecular dynamics simulations, and refinements such as hydrogen atom adjustments and coordinate corrections, resulting in more stable and reliable protein-ligand structures as well as electron cloud information. In contrast, the LBA dataset suffers from noise in the original atomic coordinates, hydrogen atom deposits, and consequently unreliable electron cloud data. These issues make it challenging for the pretrained tokenizer to capture meaningful patterns in both the electron cloud and atomic structures. We also observed similar trends during training (See Figure. 4(a)&4(b)): under identical tokenizer architectures, the model consistently achieved lower losses in both electron cloud reconstruction loss ($L^{\mathsf{ec}}$) and atomic topology/structure-related losses on MISATO, with test $L^{\mathsf{ec}}$ values being as low as a quarter of those observed on LBA. To further investigate, we separately establish two more variants of ECBind, which are fine-tuned only with Full-Atom Tokenizer (Sec. 3.2) and the Electron-Cloud Tokenizer (Sec. 3.1) on these two tasks. From Figure. 4(c), for MISATO, the two codebooks exhibited a synergistic effect, with the Electron-Cloud codebooks contributing more performance gains compared to the Full-Atom codebooks. However, for LBA, using only the Full-Atom Tokenizer was sufficient to achieve performance comparable to ECBind-stdt. This indicates that imprecise electron-cloud information provides little additional benefit when the model can already effectively capture atomic-level structures in LBA.
Effective perception of structural difference helps. Secondly, we aim to understand why the pretrained ECBind shows such significant improvements in per-structure evaluations compared to its non-pretrained counterpart. An observation is that in a binding site, there are typically over 500 atoms from the protein and fewer than 50 from the molecule. Per-structure evaluations are conducted within the same protein binding to different molecules. We hypothesize that pretraining on both atom-level and electron-cloud-level structures enables the model to capture subtle structural differences from changes in molecules, leading to more accurate predictions. In contrast, the non-pretrained ECBind, while performing well on overall evaluations, primarily focuses on protein structures. Its strong overall performance likely stems from correctly distinguishing protein structural differences rather than capturing the intermolecular interaction patterns within the binding site. To investigate this, we select two different molecules binding to the same protein and define the ‘key interacting nodes’ as atom pairs from two biomolecules within a distance of $\leq 4\ \text{\AA}$. We extract the representations of proteins’ key nodes before passing them through the decoder. Since we employed a spherical projection codebook, the representation similarity between the selected key nodes of the two binding sites is computed using cosine similarity. As shown in Figure. 5, the atom representations after tokenization exhibit greater distinctiveness, preventing the encoder from over-smoothing them and the decoder thus from confusing them, resulting in more accurate predictions.
6 Conclusion and Limitation
We propose ECBind which integrates electronic and atomic structures for protein-ligand interaction learning. One limitation of ECBind is the additional computational cost associated with pretraining, which often relies on high-quality binding complex structures to achieve significant performance gains. Moreover, ECBind currently lacks fragment-level knowledge, and incorporating this information could further improve its predictive accuracy. Exploring these aspects will be our future work.
Broader Impacts
By introducing quantum-scale electron cloud information into molecular representation learning, ECBind paves the way for a deeper and more physically grounded understanding of biomolecular interactions. Its ability to capture fine-grained electronic features provides a new method of precise prediction of protein-ligand binding affinity, potentially accelerating drug discovery processes. Beyond drug design, ECBind’s approach could be extended to diverse applications, such as modeling enzyme catalysis, understanding allosteric regulation, or designing functional molecular materials with targeted reactivity profiles. Furthermore, the knowledge distillation component of ECBind ensures accessibility and efficiency, making it feasible to deploy advanced quantum-informed models in large-scale virtual screening and real-time therapeutic evaluation pipelines. This democratization of quantum-level insights—without prohibitive computational costs—may foster broader adoption of physically-aware AI in biology and chemistry, catalyzing new interdisciplinary research and innovation.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Abstract, Sec. 1: Line. 51-59
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Sec. 6., Sec. 5: Line. 298-302
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Appendix. B
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Appendix.  F: Table. 4 and Table. 5
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification:
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Sec. 5.1 and Appendix. D and E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Table. 1 and Table. 2, Figure. 3
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix. E. and G.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The anonymity is preserved, and there is no questioned ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Boarder Impacts after the conclusion sections.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The reference and the related hyperlink to github codebase.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should citep the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification:
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification:
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification:
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification:
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Gallery of Notations
Symbol Description
$\mathcal{B}$ Protein-ligand pair
$\mathcal{P}$ Protein atomic structures
$\mathcal{M}$ Molecule (ligand) atomic structures
$\bm{V}$ Node set of $\mathcal{P}$ and $\mathcal{M}$
$\bm{E}$ Edge set of $\mathcal{P}$ and $\mathcal{M}$
$N_{\mathrm{fa}}$ Number of atoms
$N_{\mathrm{ec}}$ Number of electron cloud signals
$\mathcal{I}_{\mathrm{rec}}$ Receptors’ index set
$\mathcal{I}_{\mathrm{lig}}$ Ligands’ index set
$a_{i}$ Atom types
$\bm{\mathrm{x}}^{\mathsf{a}}_{i}$ 3D position of atom $a_{i}$
$\mathcal{D}$$\{(u_{i},\bm{\mathrm{x}}^{\mathsf{u}}_{i})\}_{1\leq i\leq N_{\mathrm{ec}}}$, Electron cloud signals
$\bm{\mathrm{x}}^{\mathsf{u}}_{i}$ Eletron density’s position
$u_{i}$ Electron density at position $\bm{\mathrm{x}}^{\mathsf{u}}_{i}$
$\bm{z}^{\mathsf{u}}_{i}$ High-dimensional embedding extracted by $u_{i}$
$\bm{z}^{\mathsf{a}}_{i}$ High-dimensional embedding extracted by $a_{i}$
$\bm{Z}^{\mathsf{u}}_{i}$ $(\bm{z}^{\mathsf{a}}_{i},\bm{z}^{\mathsf{u}}_{j_{1}},\ldots,\bm{z}^{\mathsf{u}}_{j_{K}})$, embedding sequence of patches.
$\bm{X}^{\mathsf{u}}_{i}$ $(\bm{\mathrm{x}}^{\mathsf{a}}_{i},\bm{\mathrm{x}}^{\mathsf{u}}_{j_{1}},\ldots,\bm{\mathrm{x}}^{\mathsf{u}}_{j_{K}})$, the corresponding position sequence.
$\gamma_{\mathrm{pt}}$ Predefined smoothing coefficient
$\bm{Q}^{\mathsf{iv}}$ Invariant-feature queries
$\bm{Q}^{\mathsf{ge}}$ Geometric-feature queries
$\bm{K}^{\mathsf{iv}}$ Invariant-queries keys
$\bm{K}^{\mathsf{ge}}$ The geometric-queries keys
$\bm{V}$ Values in attention
$H$ Number of attention heads
$\bm{\mathrm{A}}^{\mathsf{iv}}$ Classical attention weight matrix
$\bm{\mathrm{A}}^{\mathsf{ge}}$ Geometric attention weight matrix
$\bm{\mathrm{A}}$ Final attention weight matrix
$\mathrm{pdist}$ Pairwise-distance function
$\bm{O}$ GeoMHA-based transformer’s output
$\phi^{\mathsf{ec}}(\cdot,\cdot)$ GeoMHA-based transformer on electron-cloud-level message passing
$\phi^{\mathsf{fa}}(\cdot,\cdot)$ Full-atom GeoMHA-based transformer
$\mathcal{A}^{\mathsf{u(1)}},\mathcal{A}^{\mathsf{u(2)}}$ Codebook for 3D electronic structures and 2D atomic attributes
$\bm{h}^{\mathsf{u}}_{i}$ Patchwise representation
$\bm{h}_{i}^{\prime\mathsf{u}}$ Discrete codes of $\bm{h}^{\mathsf{u}}_{i}$
$L^{\mathsf{ec}}$ The loss function of electron
$L^{\mathsf{at}}$ The loss function of atom
$L^{\mathsf{st}}$ The intra-molecular binned distance classification loss
$L^{\mathsf{it}}$ The inter-molecular interaction loss
$L^{\mathsf{cmt}}$ The commitment loss
$L^{\mathsf{u}}$ The training objective for the tokenizer
$\alpha^{\mathsf{u}}$ The loss weight
Table 3: Gallery of Notations
Appendix B Proof of Invariance
Lemma. For the feature $\bm{H}\in\mathbb{R}^{N\times D}$ which is invariant to the translation, rotation and reflection, then for any function $F:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D^{\prime}}$, such that $\bm{H}^{\prime}=F(\bm{H})\in\mathbb{R}^{N\times D^{\prime}}$ will be invariant to the corresponding operations.
Then the invariance of $\bm{O}=\bm{AV}$ should prove that $\bm{A}$ and $\bm{V}$ are invariance to geometric translation. In this way, since $\bm{V}$ is the representation after linear transformation of $\bm{Z}$, which is the atom type embedding, it is obvious that $\bm{V}$ is invariant. For the attention matrix $\bm{A}=\text{softmax}(\bm{A}^{\mathsf{iv}}-\bm{A}^{\mathsf{ge}})$, $\bm{A}^{\mathsf{iv}}$ is the direct product of two invariant feature, and $\bm{A}^{\mathsf{ge}}$ is to measure the distance, which in detail, for the rotation matrix $\bm{\mathrm{R}}$:
$\frac{\text{softplus}(w)}{\sqrt{3}}\mathrm{pdist}(\bm{\mathrm{R}}\bm{X}\cdot\bm{Q},\bm{\mathrm{R}}\bm{X}\cdot\bm{K})$ (9)
$=$ $\frac{\text{softplus}(w)}{\sqrt{3}}\mathrm{pdist}(\bm{X}\cdot\bm{Q},\bm{X}\cdot\bm{K})$
$=$ $\bm{A}^{\mathsf{ge}}$
because the pairwise distance function is invariant to rotation. As we apply the zero-center-of-mass technique, and $\bm{\mathrm{R}}$ is the rotation matrix satisfies $\bm{\mathrm{R}}\bm{\mathrm{R}}^{\mathsf{T}}=\bm{\mathrm{I}}$, the translation vector $\bm{\mathrm{t}}=\bm{0}$. Therefore, $\bm{A}^{\mathsf{ge}}$ is invariant to the rotation and translation, leading to the invariance of $\bm{A}$.
Appendix C Tokenizing Atom-level Structure
C.1 Full-atom Tokenizer
The macro-environment around the $i$-th atom is defined similarly to the patchification of electron clouds. The full-atom GeoMHA-based transformer $\phi^{\mathsf{fa}}(\cdot,\cdot)$ encode the macro-environment as
$\quad\quad\quad\bm{h}^{\mathsf{a}}_{i}$ $=[\phi^{\mathsf{fa}}(\bm{Z}^{\mathsf{a}}_{i},\bm{X}^{\mathsf{a}}_{i})]_{1},\quad$ where (10)
$\bm{Z}^{\mathsf{a}}_{i}$ $=(\bm{z}_{i}^{\mathsf{a}},\bm{z}_{i_{1}}^{\mathsf{a}},\ldots,\bm{z}_{i_{K}}^{\mathsf{a}}),\quad\quad\quad\quad$ and
$\bm{X}^{\mathsf{a}}_{i}$ $=(\bm{\mathrm{x}}_{i}^{\mathsf{a}},\bm{\mathrm{x}}_{i_{1}}^{\mathsf{a}},\ldots,\bm{\mathrm{x}}_{i_{K}}^{\mathsf{a}}).$
Here, the $i_{k}$-th atom belongs to the neighborhood of the $i$-th atom based on their Euclidean distance.
Two codebooks, $\mathcal{A}^{\mathsf{a}}=\mathcal{A}^{\mathsf{a}(1)}\times\mathcal{A}^{\mathsf{a}(2)}$, are constructed to capture the 3D conformational geometry and 2D atomic topology, following the same approach as Eq. 2. For $\bm{h}_{i}^{\prime\mathsf{a}(2)}$ obtained from $\mathcal{A}^{\mathsf{a}(2)}$, the pretraining follows the Masked AutoEncoder (MAE) tasks described in Sec. 3.1.3, leading to the loss function $L^{\mathsf{at}}$. For $\bm{h}_{i}^{\prime\mathsf{a}(1)}$, we randomly mask the positions of a subset of atoms in the input and reconstruct the geometries, as discussed in the following sections.
C.2 Structure Decoder and Loss Function.
A classical transformer is employed to decode the tokenized representation $\bm{h}_{i}^{\prime\mathsf{a}(1)}$ into a structural one, defined as $\bm{h}_{i}^{\mathsf{st}}=\psi^{\mathsf{st}}(\bm{h}_{i}^{\prime\mathsf{a}(1)})$.
Two types of loss functions are used: (i) intra-molecular binned distance classification loss and (ii) inter-molecular interaction loss.
For (i), we first compute the ground-truth Euclidean distance between atoms as $\|\bm{\mathrm{x}}_{i}^{\mathsf{a}}-\bm{\mathrm{x}}_{j}^{\mathsf{a}}\|_{2}$. This distance is divided into discrete intervals $[0,2,3,4,\dots,20,+\infty]$, resulting in a pairwise distance label $y^{\mathsf{st}}_{i,j}$. To generate pairwise logits for structure prediction, we employ a pairwise_proj_head, which processes input embeddings $\bm{H}\in\mathbb{R}^{N\times D}$ and outputs logits $\bm{O}=\text{logit}\in\mathbb{R}^{L\times L\times d^{\prime}}$. The method follows these steps:
Algorithm 1 pairwise_proj_head
0:  $\bm{H}\in\mathbb{R}^{L\times d}$
1:  $\bm{Q},\bm{K}\leftarrow\text{Linear}(\bm{H}),\text{Linear}(\bm{H})$
2:  $\text{prod}_{i,j,:},\text{diff}_{i,j,:}\leftarrow\bm{Q}_{j,:}\odot\bm{K}_{i,:},\bm{Q}_{j,:}-\bm{K}_{i,:}$
3:  $\bm{O}\leftarrow\text{MLP}([\text{prod};\text{diff}])\in\mathbb{R}^{L\times L\times d^{\prime}}$
3:  $\bm{O}$
Therefore, by employing the discussed pair-wise head, we map the pair representations $(\bm{h}_{i}^{\mathsf{st}},\bm{h}_{j}^{\mathsf{st}})$ to logits $\mathrm{logit}_{i,j}^{\mathsf{st}}\in\mathbb{R}^{20}$. Cross-entropy is applied for $i,j\in\mathcal{I}_{\mathrm{rec}}\text{ or }i,j\in\mathcal{I}_{\mathrm{lig}}$ as the loss function:
$L^{\mathsf{st}}=\mathrm{CE}_{i,j}(\mathrm{logit}_{i,j}^{\mathsf{st}},y^{\mathsf{st}}_{i,j}).$ (11)
For (ii), a binary interaction label $y^{\mathsf{it}}_{i,j}$ is defined, indicating whether two atoms are interacting based on their distance ( $\leq$ 6Å). Another pairwise head maps $(\bm{h}_{i}^{\mathsf{st}},\bm{h}_{j}^{\mathsf{st}})$ to $\text{prob}_{i,j}^{\mathsf{it}}$, representing the probability of interaction. Binary cross-entropy is used as the loss function:
$L^{\mathsf{it}}=\mathrm{BCE}_{i,j}(\text{prob}_{i,j}^{\mathsf{it}},y^{\mathsf{it}}_{i,j}),$ (12)
where $i\in\mathcal{I}_{\mathrm{rec}},j\in\mathcal{I}_{\mathrm{lig}}$ or $i\in\mathcal{I}_{\mathrm{lig}},j\in\mathcal{I}_{\mathrm{rec}}$.
Appendix D Baseline Method Introduction
In this section, we provide a brief overview of the baseline methods selected for comparison in this study, as listed:
  • SchNet is a deep learning architecture designed for modeling quantum interactions in molecules using continuous-filter convolutional layers. It predicts total energy and interatomic forces while adhering to quantum-chemical principles, such as rotational invariance and smooth potential energy surfaces.
  • EGNN is a 3D equivariant graph neural network that not only maintains performance but also avoids the complex computations of higher-order representations in intermediate layers, extending equivariant properties to dimensions beyond 3D.
  • DimeNet is a graph neural network that enhances quantum mechanical property predictions by incorporating directional message passing. Unlike traditional models that use only atomic distances, DimeNet embeds directional information between atoms, ensuring rotational equivariance. It uses spherical Bessel functions and spherical harmonics for efficient, orthogonal representations, outperforming Gaussian radial basis functions while using fewer parameters.
  • GemNet is a geometric message-passing neural network that uses directed edge embeddings and two-hop message passing, making it a universal approximator for predictions that are invariant to translation and equivariant to permutation and rotation.
  • TorchMD is a framework designed for molecular simulations that combines classical and machine learning-based potentials. It represents all force computations, such as bond, angle, and Coulomb interactions, using PyTorch arrays and operations, enabling seamless integration with neural network potentials. The framework supports learning and simulating both ab initio and coarse-grained models, and has been validated through standard Amber all-atom simulations and protein folding tasks.
  • MACE is an equivariant message passing neural network (MPNN) model that utilizes higher-body order messages, specifically four-body messages, which reduces the number of required message passing iterations to just two, making it fast and highly parallelizable.
  • Equiformer is a graph neural network that extends Transformer architectures to 3D atomistic graphs by incorporating SE(3)/E(3)-equivariant features using irreducible representations (irreps). It improves the original Transformer model by replacing standard operations with their equivariant counterparts and integrating tensor products, enabling efficient encoding of equivariant information. Additionally, Equiformer introduces a novel equivariant graph attention mechanism, which enhances performance by using multi-layer perceptron attention and non-linear message passing.
  • LEFTNet is a geometric GNN that incorporates local substructure encoding (LSE) and frame transition encoding (FTE) to effectively represent global geometric information from local patches, achieving state-of-the-art performance on molecular property prediction tasks.
  • HEGNN is a high-degree version of EGNN to increase the expressivity by incorporating high-degree steerable vectors while still maintaining EGNN’s advantage through the scalarization trick.
Appendix E Training Detail
Experiments are conducted based on Pytorch 2.4.2 on a hardware platform with Intel(R) Xeon(R) Gold 6240R @ 2.40GHz CPU and NVIDIA A100 GPU. We give the detailed parameters for training the included models in Table. 5,  6 and  7. Besides, the training epoch for LBA and LEP, MISATO is 20, while the finetuning for LBA is 10 and for MISATO and LEP is 20.
Appendix F Ablation Study
We here aim to figure out the effects of hyper-parameters, including (i) codebook size, which is $|\mathcal{A^{\mathsf{u(1)}}}|=|\mathcal{A^{\mathsf{u(2)}}}|$ , (ii) embedding size, which is $D$, (iii) layer number of encoder $N_{\mathrm{enc}}$, (iv) layer number of decoder $N_{\mathrm{dec}}$,
Table 4: The ablation study. PS means per-structure.
D Pearson_PS Spearman_PS
32 0.226 0.209
64 0.314 0.297
128 0.423 0.448
256 0.431 0.445
codebook_size Pearson_PS Spearman_PS
256 0.378 0.366
512 0.389 0.392
1024 0.454 0.433
2048 0.431 0.445
n_layer_enc Pearson_PS Spearman_PS
32 0.226 0.209
64 0.314 0.297
128 0.423 0.448
256 0.431 0.445
n_layer_dec Pearson_PS Spearman_PS
1 0.245 0.248
2 0.287 0.305
3 0.431 0.445
6 0.392 0.410
Table 5: Hyperparameters for each baseline on each task.(Part 1)
hyperparameter MISATO LBA LEP hyperparameter MISATO LBA LEP
SchNet-fragment
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
SchNet-atom
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
SchNet-hierarchical
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
DimeNet++-fragement
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
DimeNet++-atom
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
DimeNet++-hierarchical
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
EGNN-fragment
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
EGNN-atom
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
EGNN-hierarchical
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
ET-fragment
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
ET-atom
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
ET-hierarchical
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
GemNet-fragment
$d_{h}$ 64 64 128 max_n_vertex 1500 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
GemNet-atom
$d_{h}$ 32 64 128 max_n_vertex 200 500 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
GemNet-hierarchical
$d_{h}$ 32 64 128 max_n_vertex 200 500 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
Table 6: Hyperparameters for each baseline on each task.(Part 2)
hyperparameter MISATO LBA LEP hyperparameter MISATO LBA LEP
MACE-fragment
$d_{h}$ 64 64 128 max_n_vertex 1000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
MACE-atom
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
MACE-hierarchical
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
Equiformer-fragment
$d_{h}$ 32 64 128 max_n_vertex 700 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
Equiformer-atom
$d_{h}$ max_n_vertex - - -
lr - - - final_lr - - -
max_epoch - - - save_topk - - -
Equiformer-hierarchical
$d_{h}$ max_n_vertex - - -
lr - - - final_lr - - -
max_epoch - - - save_topk - - -
LEFTNet-fragment
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
LEFTNet-atom
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
LEFTNet-hierarchical
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
GET-fragment
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
GET-atom
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 1500
lr $10^{-3}$ $10^{-3}$ $5\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 90 save_topk 3 3 7
GET-hierarchical
$d_{h}$ 64 64 128 max_n_vertex 2000 2000 500
lr $10^{-3}$ $10^{-3}$ $10^{-3}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-4}$
max_epoch 10 10 70 save_topk 3 3 5
Table 7: Hyperparameters for our ECBind on each task.
hyperparameter MISATO LBA LEP hyperparameter MISATO LBA LEP
ECBind-nptn
$D$ 256 256 256 $d_{r}$ 16 32 64
lr $3\times 10^{-5}$$\times 10^{-4}$ $3\times 10^{-4}$ final_lr $10^{-6}$ $10^{-6}$ $10^{-6}$
max_epoch 20 20 20 save_topk 3 3 3
n_layers_enc 3 3 3 n_layers_dec 3 3 3
codebook_size 2048 2048 2048
Appendix G Time and Memory Analysis
We report the average inference time on the total test set and peak memory usage in the training of all models. The results are averaged over 3 independent runs using a single A100 GPU. Since our models are totally transformer-based which can integrate linear-attention mechanisms like FlashAttention, it is extremely fast in inference and much faster than graph-based models like EGNN and GemNet.
Table 8: Comparison of inference time and memory consumption across models on MISATO.
Model Time (s) Memory (MB)
EGNN-fragment 14.72 361.62
EGNN-atom 20.24 2211.86
EGNN-hierarchical 27.74 2209.14
SchNet-fragment 14.77 819.17
SchNet-atom 17.57 2982.90
SchNet-hierarchical 23.33 2861.80
TorchMD-fragment 20.13 897.87
TorchMD-atom 24.21 5838.56
TorchMD-hierarchical 33.92 5835.98
LEFTNet-fragment 24.25 1573.83
LEFTNet-atom 32.86 13592.21
LEFTNet-hierarchical 32.97 13592.21
GemNet-fragment 69.02 20041.35
GemNet-atom 448.26 15087.50
GemNet-hierarchical 633.70 13972.60
DimeNet-fragment 53.95 5371.92
DimeNet-atom 77.98 29456.64
DimeNet-hierarchical 80.89 27150.04
Equiformer-fragment 123.96 34447.65
Equiformer-atom 873.22 17780.71
Equiformer-hierarchical 881.35 17970.31
MACE-fragment 59.85 2243.75
MACE-atom 91.36 23564.17
MACE-hierarchical 85.24 23709.81
GET-fragment 59.08 4523.77
GET-atom 61.68 5361.85
GET-hierarchical 118.14 6691.07
HEGNN-fragment 27.93 643.68
HEGNN-atom 32.27 2298.35
HEGNN-hierarchical 41.56 2300.65
ECBind-nptn 9.13 6724.58
ECBind-stdt 9.16 6724.58
ECBind-ptn 11.48 9302.74